# UFM: A Simple Path towards Unified Dense Correspondence with Flow

### v1

@@ 2 Related Works
Dense wide-baseline matchers suppress their sparse counterparts since DKM [14], which first obtains a robust, coarse match from patch features and uses regressive warp-refiners to upsample the prediction resolution. RoMa [15] builds upon DKM by using a frozen image foundation model (DINOv2 [41]) for its coarse matching and uses separate convolution-based encoders to provide fine details to warp-refiners. Despite being robust and accurate, both methods have a heavy architecture that limits their application to compute-limited scenarios. We show that our method can achieve similar robustness and accuracy while being about 6×6$\times$6 × faster.
@@
Recent advancements in end-to-end learning have also encouraged scaling a generic architecture for correspondence. CroCoV2 [64] shows that optical flow can be directly regressed from its backbone pre-trained on the cross-image-completion task. However, they stopped at low resolution and required a sliding window method to infer at high resolution, which failed to capture correspondences across windows. Furthermore, CroCov2 doesn’t train the two-view transformer from scratch to directly regress flow. More recent follow-up MASt3R [30] finetuned DUSt3R [61] to output pixel-wise feature descriptors and proposed a fast reciprocal matching to decode sparse matches efficiently. However, this paradigm does not provide dense matches and is prohibitively slow without subsampling.
@@ 3.1.  UFM Architecture
Given two images $I_{1},I_{2}\in\mathbb{R}^{3\times H\times W}$ as input, our Unified Flow and Matching (UFM) model (Fig. 2) predicts the visually grounded dense correspondences and covisibility:
@@
To achieve this, UFM employs a simple end-to-end transformer with multiple benefits in modeling power for large displacements and speed: (1) No pixel is left behind. Unlike the commonly used coarse-to-fine paradigm [63, 68, 15], which restricts attention to local regions in the cost volume and assumes uniform motion within patches, transformer-based models estimate flow features with a global receptive field. This prevents uncorrectable motion features, which may be dominant due to motion patterns within the patch. In coarse-to-fine methods, such errors are hard to correct later, as attention becomes increasingly localized at finer resolutions. This effect is evident in Fig. 4, where coarse-to-fine flow results in pixelated artifacts near inclined thin structures. (2) Operating at the patch level when constructing flow is fast, and DPT [43] enables detailed decoding. (3) Structural simplicity enables easy optimization and potential for additional simple fine-to-fine refinements without a huge impact on efficiency. We elaborate on the end-to-end transformer further below.
@@
While we find that the regression of dense correspondence (flow) is robust, it is not always precise (e.g., see average EPE & outlier numbers for UFM 560 in Section 3.4). Hence, we designed a simple classification-based local refinement technique to improve the accuracy of UFM’s inlier predictions. We take inspiration from MASt3R [30]’s design to regress pixel-wise matching features based on transformer backbone features. Additionally, to capture fine details for the refinement, we added a U-Net encoder following RoMa [15]. As shown in Fig. 3, we differ from MASt3R [30] in how we leverage the refinement features for correspondence: as opposed to matching dense features across the entire image (global search), we use the regressed flow from UFM’s DPT to guide the feature matching around a small $7\times 7$ neighborhood, thereby leading to 60×60$\times$60 × efficiency over MASt3R (Section 3.4). In particular, we compute the attention between each pixel and its local $7\times 7$ neighborhood determined by the regressed flow and use the weighted sum of coordinates by the softmax attention as the residual to update the initially regressed flow.
@@ 3.2.  Training Objective
We trained with a robust regression loss [1], following the approach in RoMa [15], which focuses its gradient on inlier predictions with small errors—typically around 1 to 2 pixels. We selected this loss for two main reasons. First, it encourages precise learning from reliable matches by emphasizing small residuals. Second, it reduces the impact of incorrect data during training, as robust losses exhibit vanishing gradients for large flow errors, which are commonly caused by unmatchable pairs. Specifically, we used the generalized Charbonnier loss with parameters $\alpha=0.5$ and $c=0.24$.
@@ 3.3.  Combining Flow and Matching Datasets
Monkaa [34] $8.6$ k $8.6$ k $24$ Synthetic $5$ k
HD1K [37] $1081$ $1046$ $35$ Real $5$ k
We compiled a unified dataset consisting of 12 datasets spanning diverse sources, motion patterns, and environments from both wide-baseline matching and optical flow domains, as detailed in Tab. 3.3. The collection of these datasets features diverse indoor, outdoor, in-the-wild, and dynamic scenes.
@@ 3.4.  Training Details
We train the network with a longest side resolution of $560$ (with aspect ratios varying from 3:1 to 1:1) for $48$ epochs with data as specified in Section 3.3. All our datasets permit academic research, and the publicly released UFM model weights will be licensed following this. We use a peak learning rate of $1\cdot 10^{-4}$ for the global attention transformer and DPT heads and $5\cdot 10^{-6}$ for the encoder to preserve DINOv2 pre-training. This contrasts with the frozen DINOv2 used by RoMA [15] (which we find suboptimal), and we provide further insights in the supplementary. We use AdamW optimizer with a cosine decay learning rate schedule using $10\%$ linear warmup, $0.05$ weight decay, and $\beta=\{0.9,0.95\}$. Since most of our data has bidirectional correspondence, we symmetrize the batches. This leads to an effective batch size of $96$ pairs, where half of them are unique. The training takes 4 days on $8$ H100 GPUs. We name this checkpoint as UFM 560.
Some downstream tasks, like visual odometry [42], require sub-pixel accuracy, making high-resolution images essential. However, training at high resolution is computationally expensive. To address this, we bootstrap a high-resolution model, UFM 980, from UFM 560. The wide-baseline datasets do not have depth annotations at a high resolution (1K), and upsampling the pre-computed flow at lower resolutions would be sub-optimal for sub-pixel training. Hence, we train with 10×10$\times$10 × lower learning rates than the 560 training on all optical flow data for 15 epochs. Furthermore, we change the supervision range to all pixels to follow the standard evaluation protocol in optical flow.
Table 2: Wide Baseline Dense Correspondence: Zero-shot dense correspondence evaluation at all covisible pixels. We report the AEPE and outlier rates at thresholds of 1, 2, and 5 pixels. UFM outperforms all dense methods by a large margin and matches MASt3R’s performance, despite MASt3R’s advantage in selecting its confident pixels, while being 60×60$\times$60 × faster.
@@
[…]
@@ 4.1.  Zero-Shot Wide-Baseline Correspondence
In Section 3.4, despite giving MASt3R an advantage, we showcase that UFM significantly outperforms all dense methods in precision, while achieving nearly 60×60$\times$60 × lower runtime than MASt3R — the only method with comparable precision. Furthermore, UFM significantly outperforms all dense methods, achieving on average $62\%$ less EPE and 6.7×6.7$\times$6.7 × better runtime compared to the best dense baseline, RoMa.
@@
We follow DKM [14] and evaluate UFM for pose estimation. As shown in Table 3, UFM achieves the best accuracy on ETH3D and TA-WB benchmark, and second place on Scannet-1500 (despite not being trained on this dataset). This performance shows that UFM’s correspondence is well-balanced and suitable for 3D geometric tasks. Moreover, we observe a notable improvement by adding refinement on top of UFM, highlighting the potential for integrating other refinement techniques on top of the base model for further improvement.
@@ 4.2.  Optical Flow Correspondence
We compare our approach to all optical flow methods in Section 4.1 and RoMa, using the checkpoint trained on FlyingChairs, FlyingThings, and TartanAir (SEA-RAFT only)—i.e., the best trained model before violating the zero-shot setting.
@@ Results
Section 4.1 shows that UFM, without any refinement, achieves state-of-the-art zero-shot performance on Sintel-Final and KITTI in terms of both EPE and most pixel outlier metrics, while also delivering competitive performance on Sintel-Clean. These results demonstrate that the UFM base model has strong generalization and precision to be combined with existing refinement techniques.
@@ 4.3.  Generalizable Matching on Ego-Exo 4D
We ran UFM on images from the Ego-Exo4D [16], which features videos captured in first and third person view across diverse scenes. As shown in Fig. 4, compared to RoMA, UFM achieves strong generalization & robust matching.
@@ 4.4.  Insights towards Unified Correspondence
We conduct an ablation study to see if UFM benefits from unified training on merged data as opposed to training on specialized data only. Specifically, we train UFM only on optical, wide baseline, and the combination for $100+20$ epochs at $224$ $560$ resolutions. Across the different variants, UFM sees each data point the same number of times. Although the total number of gradient steps differs, the number of epochs is large enough for the training to have effectively converged. We then evaluate optical flow and dense wide-baseline performance as in the previous sections.
@@
In Table 5, UFM outperforms it’s own specialized variants, thereby indicating a mutual improvement when merging the two data types. For optical flow, we observe that adding wide-baseline data brings a $20\%-80\%$ decrease in EPE, especially on the KITTI dataset. For wide-baseline, we observe that adding optical flow data brings a $3.2\%$ relative decrease in 1, 2 pixel outlier rates.
@@
To test the scalability of existing architectures, we trained SEA-RAFT, UniMatch, RoMa, and UFM on the same unified data (Section 3.3). Each architecture is trained with its original loss functions as specified in the respective papers. We recorded the validation set EPE at data bound ($35$ epochs, $22.5$ M pairs) and compute bound ($32$ hours on $8$ H100 GPU) to measure scalability. Fig. 5 shows that UFM performs best on all datasets at both data and compute bound. This indicates the benefits of using a simple architecture to scale on large amounts of data, where UFM shows significantly increasing performance with compute on harder datasets like ScanNet++ and TartanAir.
@@ 5 Limitations
While UFM represents an exciting development in constructing models for unified dense image correspondence, some limitations remain with semantic matching capabilities. As shown in Fig. 6, on WxBS [39], we find that UFM works on challenging image pairs that demonstrate scale, viewpoint, texture, and illumination and tends to struggle with extreme seasonal changes and matching across spectrums, i.e., visual to very dark infrared (thermal). RoMA [15] is robust to such semantic changes due to the coarse patch correlation provided by frozen DINOv2 [41] features, with the help of additional fine features from ConvNet in its upsampling process. We find that freezing the encoder does not benefit an end-to-end transformer architecture such as UFM. As shown in Appendix E, we find that freezing the pre-trained DINOv2 image encoder leads to a significant drop in dense correspondence performance. Opportunities may lie in complementing frozen DINOv2 features from other sources. Furthermore, although we constrain the learning rate to remain relatively small to preserve DINOv2’s pretraining, we find that DINOv2 can still deviate significantly during extensive training and lose some of its semantic matching abilities. Through preliminary exploration, we find that this can be mitigated by incorporating semantic matching data, semantic preservation losses, or specialized fine-tuning that limits the extent of deviation from the pre-trained weights. We aim to address this in future releases of UFM.
@@ 6 Conclusion
Figure 6: WxBS Benchmarking [39]: We find that UFM: (a) outperforms MASt3R [30], another end-to-end transformer trained on large-scale data for correspondence; (b-f) performs well on images with scale, viewpoint, illumination, and seasonal changes, and (g-i) struggles with pairs showing extreme coupled season, illumination, and scale changes or captured across different imaging spectrums, where RoMA [15] is more robust. We provide further insights in Section 5 and believe the primary reason to be the preservation of semantic matching capabilities in the pre-trained image encoder.
@@ Appendix A Computing Covisibility Mask
Computing the covisibility mask for all datasets in Section 3.3 is a key step to support unified training. In this section, we detail the exact protocol and parameters we used to compute the covisibility mask for all datasets, summarized in Table S.1. We will begin with a general principle of using depth reprojection error to compute covisibility, and then detail its application to three data categories: (1) Static Scenes, (2) Optical Flow, and (3) Rigid Posed Objects.
@@ A.1.  Covisibility from Depth Reprojection Error
All dataset categories use the same covisibility thresholding scheme, with dataset-specific parameters summarized in Table S.1. However, each category differs in how the 3D points $p_{s}$ and $p_{t}$, and ultimately the error $e$, are computed. We describe these procedures in the following subsections.
@@ A.2.  Computing Correspondence and Covisibility
Formally, we use uniform camera projection model $\pi$ and all information as described in static scenes, optical flow ground truth $\phi^{gt}\in\mathbb{R}^{2\times H\times W}$, and depth (disparity) change $D_{1\to 2}\in\mathbb{R}^{H\times W}$. As optical flow describes how a source pixel $i_{s}$ moves to the target pixel $i_{t}$ in the image space, depth change details how the underlying 3D point changes in its depth. Specifically, the 3D point refered by $i_{s}$ at the source image with depth $D_{1}(i_{s})$ would move to pixel $i_{t}=i_{s}+\phi^{gt}(i_{s})$ in the target image, with an updated depth of $D_{1}(i_{s})+D_{1\to 2}(i_{s})$. Thus, we can compute the source point in the target time and the projection error (similar to Eq. S.3) as:

### v2

@@ 2 Related Works
Dense wide-baseline matchers suppress their sparse counterparts since DKM [14], which first obtains a robust, coarse match from patch features and uses regressive warp-refiners to upsample the prediction resolution. RoMa [15] builds upon DKM by using a frozen image foundation model (DINOv2 [46]) for its coarse matching and uses separate convolution-based encoders to provide fine details to warp-refiners. Despite being robust and accurate, both methods have a heavy architecture that limits their application to compute-limited scenarios. We show that our method can achieve similar robustness and accuracy while being about $6\times$ faster.
@@
Recent advancements in end-to-end learning have also encouraged scaling a generic architecture for 3D reconstruction and correspondence [67, 45, 65, 31, 29, 26]. Perceiver IO [24] shows its architecture can solve diverse vision tasks, including optical flow. CroCoV2 [70] also shows that optical flow can be directly regressed from its backbone pre-trained on the cross-image-completion task. However, CrocoV2 stopped at low resolution and both methods required a sliding window method to infer at high resolution, which failed to capture correspondences across windows. Furthermore, CroCov2 doesn’t train the two-view transformer from scratch to directly regress flow. More recent follow-up MASt3R [34] finetuned DUSt3R [67] to output pixel-wise feature descriptors and proposed a fast reciprocal matching to decode sparse matches efficiently. However, this paradigm does not provide dense matches and is prohibitively slow without subsampling.
@@ 3.1. UFM Architecture
Given two images $I_{1},I_{2}\in\mathbb{R}^{3\times H\times W}$ as input, our Unified Flow and Matching (UFM) model (Fig.˜2) predicts the visually grounded dense correspondences and covisibility:
@@
In particular, UFM employs a simple end-to-end transformer with a CNN head to directly regress the flow value, unlike prior methods [69, 74, 15] that first construct an initial solution at the patch level and then progressively upsample to the pixel level. We empirically find the paradigm used in prior methods to yield suboptimal accuracy, as shown in Appendix˜F. Beyond accuracy, direct flow regression offers additional advantages: (1) Performing attention at the patch level is fast for flow prediction, while the DPT [48] enables predictions directly at the pixel level. (2) Structural simplicity enables easy optimization and potential for additional simple pixel-level refinements without a huge impact on efficiency. We elaborate on the end-to-end transformer further below.
@@
While we find that the regression of dense correspondence (flow) is robust, it is not always precise (e.g., see average EPE & outlier numbers for UFM 560 in Table˜2). Hence, we designed a simple classification-based local refinement technique to improve the accuracy of UFM’s inlier predictions. We take inspiration from MASt3R [34]’s design to regress pixel-wise matching features based on transformer backbone features. Additionally, to capture fine details for the refinement, we added a U-Net encoder following RoMa [15]. As shown in Fig.˜3, we differ from MASt3R [34] in how we leverage the refinement features for correspondence: as opposed to matching dense features across the entire image (global search), we use the regressed flow from UFM’s DPT to guide the feature matching around a small $7\times 7$ neighborhood, thereby leading to $60\times$ efficiency over MASt3R (Table˜2), or about $24\times$ considering the follow-up acceleration Speedy MASt3R [35]. In particular, we compute the attention between each pixel and its local $7\times 7$ neighborhood determined by the regressed flow and use the weighted sum of coordinates by the softmax attention as the residual to update the initially regressed flow.
@@ 3.2. Training Objective
We trained with a robust regression loss [1], following the approach in RoMa [15], which focuses its gradient on inlier predictions with small errors—typically around 1 to 2 pixels. We selected this loss for two main reasons. First, it encourages precise learning from reliable matches by emphasizing small residuals. Second, it reduces the impact of incorrect data during training, as robust losses exhibit vanishing gradients for large flow errors, which are commonly caused by unmatchable pairs. Specifically, we used the generalized Charbonnier loss with parameters $\alpha=0.5$ and $c=0.24$. See Appendix˜H for more visualization and discussion.
@@ 3.3. Combining Flow and Matching Datasets
Monkaa [39] $8.6$ k $8.6$ k $24$ Synthetic 55 k
HD1K [42] $1081$ $1046$ $35$ Real 55 k
We compiled a unified dataset consisting of 12 datasets spanning diverse sources, motion patterns, and environments from both wide-baseline matching and optical flow domains, as detailed in Tab. 1. The collection of these datasets features diverse indoor, outdoor, in-the-wild, and dynamic scenes.
@@ 3.4. Training Details
We train the network with a longest side resolution of $560$ (with aspect ratios varying from 3:1 to 1:1) for $48$ epochs with data as specified in Table˜1. All our datasets permit academic research, and the publicly released UFM model weights will be licensed following this. We use a peak learning rate of $1\cdot 10^{-4}$ for the global attention transformer and DPT heads and $5\cdot 10^{-6}$ for the encoder to preserve DINOv2 pre-training. This contrasts with the frozen DINOv2 used by RoMA [15] (which we find suboptimal), and we provide further insights in the supplementary. We use AdamW optimizer with a cosine decay learning rate schedule using $10\%$ linear warmup, $0.05$ weight decay, and $\beta=\{0.9,0.95\}$. Since most of our data has bidirectional correspondence, we symmetrize the batches. This leads to an effective batch size of $96$ pairs, where half of them are unique. The training takes 4 days on $8$ H100 GPUs. We name this checkpoint as UFM 560.
Some downstream tasks, like visual odometry [47], require sub-pixel accuracy, making high-resolution images essential. However, training at high resolution is computationally expensive. To address this, we bootstrap a high-resolution model, UFM 980, from UFM 560. The wide-baseline datasets do not have depth annotations at a high resolution (1K), and upsampling the pre-computed flow at lower resolutions would be sub-optimal for sub-pixel training. Hence, we train with $10\times$ lower learning rates than the 560 training on all optical flow data for 15 epochs. Furthermore, we change the supervision range to all pixels to follow the standard evaluation protocol in optical flow.
Table 2: Wide Baseline Dense Correspondence: Zero-shot dense correspondence evaluation at all covisible pixels. We report the AEPE and outlier rates at thresholds of 1, 2, and 5 pixels. UFM outperforms all dense methods by a large margin and matches MASt3R’s performance, despite MASt3R’s advantage in selecting its confident pixels, while being $60\times$ faster ($24\times$ considering follow up [35]).
@@
$\text{AUC }@\text{ }5^{\circ}$ $\uparrow$ @ $\text{ }10^{\circ} \uparrow @$ $\text{ }15^{\circ} \uparrow$ $\text{AUC }@\text{ }5^{\circ}$ $\uparrow$ @ $\text{ }10^{\circ} \uparrow @$ $\text{ }15^{\circ} \uparrow$ $\text{AUC }@\text{ }10^{\circ}$ $\uparrow$ @ $\text{ }20^{\circ} \uparrow @$ $\text{ }30^{\circ} \uparrow$
@@ 4.1. Zero-Shot Wide-Baseline Correspondence
In Table˜2, despite giving MASt3R an advantage, we showcase that UFM significantly outperforms all dense methods in precision, while achieving nearly $60\times$ lower runtime ($24\times$ considering follow up [35]) than MASt3R — the only method with comparable precision. Furthermore, UFM significantly outperforms all dense methods, achieving on average $62\%$ less EPE and $6.7\times$ better runtime compared to the best dense baseline, RoMa.
@@
We follow DKM [14] and evaluate UFM for pose estimation. As shown in Table˜3, UFM achieves the best accuracy on ETH3D and TA-WB benchmark, and second place on Scannet-1500 (despite not being trained on this dataset). This performance shows that UFM’s correspondence is well-balanced and suitable for 3D geometric tasks. Moreover, we observe a notable improvement by adding refinement on top of UFM, highlighting the potential for integrating other refinement techniques on top of the base model for further improvement.
@@ 4.2. Optical Flow Correspondence
We compare our approach to all optical flow methods in Section˜4.1 and RoMa, using the checkpoint trained on FlyingChairs, FlyingThings, and TartanAir (SEA-RAFT only)—i.e., the best trained model before violating the zero-shot setting.
@@ Results
Table˜4 shows that UFM, without any refinement, achieves state-of-the-art zero-shot performance on Sintel-Final and KITTI in terms of both EPE and most pixel outlier metrics, while also delivering competitive performance on Sintel-Clean. These results demonstrate that the UFM base model has strong generalization and precision to be combined with existing refinement techniques.
@@ 4.3. Generalizable Matching on Ego-Exo 4D
We ran UFM on images from the Ego-Exo4D [16], which features videos captured in first and third person view across diverse scenes. As shown in Fig.˜4, compared to RoMA, UFM achieves strong generalization & robust matching.
@@ 4.4. Insights towards Unified Correspondence
We conduct an ablation study to see if UFM benefits from unified training on merged data as opposed to training on specialized data only. Specifically, we train UFM only on optical, wide baseline, and the combination for $100+20$ epochs at $224$ & $560$ resolutions. Across the different variants, UFM sees each data point the same number of times. Although the total number of gradient steps differs, the number of epochs is large enough for the training to have effectively converged. We then evaluate optical flow and dense wide-baseline performance as in the previous sections.
@@
In Table˜5, UFM outperforms it’s own specialized variants, thereby indicating a mutual improvement when merging the two data types. For optical flow, we observe that adding wide-baseline data brings a $20\%-80\%$ decrease in EPE, especially on the KITTI dataset. For wide-baseline, we observe that adding optical flow data brings a $3.2\%$ relative decrease in 1, 2 pixel outlier rates.
@@
To test the scalability of existing architectures, we trained SEA-RAFT, UniMatch, RoMa, and UFM on the same unified data (Table˜1). Each architecture is trained with its original loss functions as specified in the respective papers. We recorded the validation set EPE at data bound ($35$ epochs, $22.5$ M pairs) and compute bound ($32$ hours on $8$ H100 GPU) to measure scalability. Fig. 5 shows that UFM performs best on all datasets at both data and compute bound. This indicates the benefits of using a simple architecture to scale on large amounts of data, where UFM shows significantly increasing performance with compute on harder datasets like ScanNet++ and TartanAir. We believe UFM’s effectiveness in scaling to the unified dataset is crucial for achieving mutual benefit across opticalflow and wide-baseline matching.
@@ 5 Limitations
While UFM represents an exciting development in constructing models for unified dense image correspondence, some limitations remain with semantic matching capabilities. As shown in Fig.˜6, on WxBS [44], we find that UFM works on challenging image pairs that demonstrate scale, viewpoint, texture, and illumination and tends to struggle with extreme seasonal changes and matching across spectrums, i.e., visual to very dark infrared (thermal). RoMA [15] is robust to such semantic changes due to the coarse patch correlation provided by frozen DINOv2 [46] features, with the help of additional fine features from ConvNet in its upsampling process. We find that freezing the encoder does not benefit an end-to-end transformer architecture such as UFM. As shown in Appendix˜E, we find that freezing the pre-trained DINOv2 image encoder leads to a significant drop in dense correspondence performance. Opportunities may lie in complementing frozen DINOv2 features from other sources. Furthermore, although we constrain the learning rate to remain relatively small to preserve DINOv2’s pretraining, we find that DINOv2 can still deviate significantly during extensive training and lose some of its semantic matching abilities. Through preliminary exploration, we find that this can be mitigated by incorporating semantic matching data, semantic preservation losses, or specialized fine-tuning that limits the extent of deviation from the pre-trained weights. We aim to address this in future releases of UFM.
We also acknowledge that although our training set and the TA-WB evaluation is geometrically challenging, they are biased toward static objects, potentially reducing the accuracy on dynamic ones, as shown in Appendix˜G. Future work may include more sceneflow datasets, such as ParallelDomain-4D [63], to balance the distribution.
@@ 6 Conclusion
Figure 6: WxBS Benchmarking [44]: We find that UFM: (a) outperforms MASt3R [34], another end-to-end transformer trained on large-scale data for correspondence; (b-f) performs well on images with scale, viewpoint, illumination, and seasonal changes, and (g-i) struggles with pairs showing extreme coupled season, illumination, and scale changes or captured across different imaging spectrums, where RoMA [15] is more robust. We provide further insights in Section˜5 and believe the primary reason to be the preservation of semantic matching capabilities in the pre-trained image encoder.
@@ Appendix A Computing Covisibility Mask
Computing the covisibility mask for all datasets in Table˜1 is a key step to support unified training. In this section, we detail the exact protocol and parameters we used to compute the covisibility mask for all datasets, summarized in Table˜S.1. We will begin with a general principle of using depth reprojection error to compute covisibility, and then detail its application to three data categories: (1) Static Scenes, (2) Optical Flow, and (3) Rigid Posed Objects.
@@ A.1. Covisibility from Depth Reprojection Error
All dataset categories use the same covisibility thresholding scheme, with dataset-specific parameters summarized in Table˜S.1. They only differ in how the 3D points $p_{s}$ and $p_{t}$, and ultimately the error $e$, are computed. We describe these procedures in the following subsections.
@@ A.2. Computing Correspondence and Covisibility
Formally, we use uniform camera projection model $\pi$ and all information as described in static scenes, optical flow ground truth $\phi^{gt}\in\mathbb{R}^{2\times H\times W}$, and depth (disparity) change $D_{1\to 2}\in\mathbb{R}^{H\times W}$. As optical flow describes how a source pixel $i_{s}$ moves to the target pixel $i_{t}$ in the image space, depth change details how the underlying 3D point changes in its depth. Specifically, the 3D point refered by $i_{s}$ at the source image with depth $D_{1}(i_{s})$ would move to pixel $i_{t}=i_{s}+\phi^{gt}(i_{s})$ in the target image, with an updated depth of $D_{1}(i_{s})+D_{1\to 2}(i_{s})$. Thus, we can compute the source point in the target time and the projection error (similar to Eq.˜S.3) as: